OpenStudy (anonymous):

What is 0.12 expressed as the quotient of two integers in simplest form?

jimthompson5910 (jim_thompson5910):

0.12 = twelve hundredths 0.12 = 12/100 0.12 = ??? reduce 12/100 fully to get your final answer

OpenStudy (anonymous):

3/25

jimthompson5910 (jim_thompson5910):

yep, 12/100 = 3/25

OpenStudy (anonymous):

@AngieLovesYouxx can you please pm me back